(log x base 2)^2 – log x^3 base 2 = 4, how can i solve this equation?

Grade:10

2 Answers

Arun
6 years ago
Dear student
 
Let log x base 2 = t
 
Then
t² -3t = 4
t² -3t -4 = 0
(t -4)(t +1) = 0
t = 4 and t = -1
 
Now
Log x base 2 = 4
x = 2⁴ = 16
 
Now
Log x base 2 = -1
x = 2^-1 = ½
 
 
Regards
Arun (askIITians forum expert)
ankush rathee
6 years ago
Let (log 2 x) be a So a^2-3a-4=0So (a-4)(a+1)=0So a=4,a=-1So log 2 x=4So x=2^4=16So log 2 x=-1So 2^-1=XX=2^-1=1/2